OpenStudy (anonymous):

If after 15 years the captains age will be twice the age of the first officer , if the age of the first officer after 5 years will be 25 years whats the age of the captain now ? plz explain the answers

OpenStudy (anonymous):

55yrs

OpenStudy (anonymous):

your right but can you explain plz

OpenStudy (anonymous):

age of first off. is 25 -5 =20 after 15 yrs age of first off.=20+15=35 the age of captain after 15 years=35*2=70 so his present age =70-15=55

OpenStudy (anonymous):

plz reply to my ques

OpenStudy (anonymous):

thanks alot i really do appreciate it and i will try to answer yours :)

OpenStudy (anonymous):

ok]